OpenStudy (anonymous):

6 over quantity 5 minus square root of 3

OpenStudy (mathstudent55):

\[\frac{ 6 }{ 5 - \sqrt{3} }\] You mean that?

OpenStudy (anonymous):

yes thats what i mean

OpenStudy (mathstudent55):

since you have 5 minus sq root 3, multiply top & bottom by 5 + sq root 3

OpenStudy (mathstudent55):

\[\frac{ 6 }{ 5 - \sqrt{3} } \times \frac{ 5 + \sqrt{3} }{ 5 + \sqrt{3} }\]

OpenStudy (mathstudent55):

\[\frac{ 6(5 + \sqrt{3}) }{ 5^{2} - (\sqrt{3})^{2} }\]

OpenStudy (mathstudent55):

\[\frac{ 6(5 + \sqrt{3}) }{ 25 - 3 }\]

OpenStudy (mathstudent55):

\[\frac{ 3(5 + \sqrt{3}) }{ 11 }\]

OpenStudy (anonymous):

that not one of my answers

OpenStudy (mathstudent55):

do they have the same denominator i do? 11

OpenStudy (anonymous):

theres only one with 11 quantity of 15 plus 3 square root of 3 all over 11

OpenStudy (mathstudent55):

That's the same as my ansewr by multiplying out the 3 on top. that's correct

OpenStudy (anonymous):

thanks